# spacing=nonfrench

### v3

@@ 4.2. Wall-Street Handover Design
This section describes Wall-Street’s HO process. First, to ensure uninterrupted connectivity, we offload the neighbor cell measurement task from the UEs to the surface (Section˜4.2.2). Second, to improve efficiency, we batch the HO for all in-vehicle users into a single operation (Section˜4.2.3). Finally, to maximize reliability, we implement a make-before-break connection, linking the UE to the new cell before disconnecting from the old cell (Section˜4.2.4). The process is divided into: initial attachment, preparation, decision, execution, and completion, as detailed in Fig.˜5 (overview) and Fig.˜7 (timeline). Our design leverages the standard 5G practice where nearby gNBs transmit their broadcast signals on different time-frequency resources. This prevents collisions and allows the serving gNB to capture a neighbor’s reflected signal without interfering with the active data link.
@@ 4.2.1. Initial Attachment
As illustrated in Fig.˜6(a), the initial attachment requires a coordinated beam sweep to align the gNB, Wall-Street (operating in single-beam transmissive mode), and the UE. Every $20$ ms period, the gNB broadcasts a $5$ ms burst containing 64 synchronization signal blocks (SSBs), which are transmitted across 8 different gNB Tx beam directions. For each SSB, Wall-Street refract them into $8$ angles, testing different paths into the vehicle. For each burst, the UE holds its Rx beam and measures the power. The UE tests a different beam direction for each burst. After four bursts (one per UE Rx direction), taking $80$ ms total, the UE reports the best alignment and connects with the gNB.
A conventional amplify-and-forward repeater requires aligning four independent beams: the gNB Tx beam $f_{g}$, the repeater Rx beam $w_{r}$, the repeater Tx beam $f_{r}$, and the UE Rx beam $w_{u}$. This process can be written as:
(2) $\underset{\text{repeater (Tx beam $f_{r}$)}\to\text{UE}}{\underbrace{\left\|w_{u}^{H}\bm{H_{r\to u}}f_{r}\right\|^{2}}}\cdot\underset{\text{gNB}\to\text{repeater (Rx beam $w_{r}$)}}{\underbrace{\left\|w_{r}^{H}\bm{H_{g\to r}}f_{g}\right\|^{2}}}$
where $\bm{H_{r\to u}}$ and $\bm{H_{g\to r}}$ are channel blocks from repeater to UE and from gNB to repeater. Searching $(f_{g},w_{r},f_{r},w_{u})$ across four independent codebooks leads to $O(n^{4})$ complexity, taking about $\sim$5 sec, which is too slow for moving vehicles.
Wall-Street fundamentally reduces this complexity. It does not have separate Rx and Tx modules with independent beams. Instead, each surface configuration applies a specific phase gradient that deterministically maps any incoming wave to an outgoing wave (generalized Snell’s law: $k_{out}=k_{in}+\nabla_{\phi}(\theta)$). While the outgoing angle is dependent on the incoming angle for any single configuration, this does not constrain the search because the alignment process performs an full sweep of all gNB beam directions and all surface states. The received power is then formulated as:
(3) $\left\|w_{u}^{H}\bm{H_{s\to u}}\Theta(\theta)\bm{H_{g\to s}}f_{g}\right\|^{2}$
where $\bm{H_{s\to u}}$ and $\bm{H_{g\to s}}$ are channel blocks from surface to UE and from gNB to surface. Here, $\theta$ is a discrete surface state, and $\Theta$ is the Wall-Street’s codebook. This leads to three independent codebooks $(f_{g},\theta,w_{u})$ with $O(n^{3})$ complexity (Wu and Zhang, 2019; Jiang et al., 2024). Also, Wall-Street’s angular reciprocity allows fast UL/DL switching once a state is found.
Figure 7. Wall-Street’s handover timeline: (1) Preparation, (2) Execution, and (3) Completion phases.
@@ 4.2.2. Handover Preparation
The serving gNB periodically triggers Wall-Street to enter dual-beam mode for neighbor cell measurement. The goal is to measure neighbor cell RSRP without interrupting data sessions, as shown in Fig.˜5(b).
Since the refractive link between the serving gNB, Wall-Street, and the UE is already established, the transmissive angle $\theta_{t}$ is known. Wall-Street then uses its pre-computed codebook indexed by $(\theta_{t},\theta_{r},\alpha)$ to cycle through 8 reflective angles $\theta_{r}$, and neighbor gNBs transmit SSBs in their standard beam patterns. Since the serving gNB already knows its direction toward Wall-Street, it measures these reflected SSBs on a separate channel to prevent interference with the active data link. The power split $\alpha$ allocates less power to maintain the data connection while assigning the remaining power to the reflective link. As illustrated in Fig.˜6(b), this search takes one 5 ms burst. This allows the serving gNB to measure the neighbor RSRPs on behalf of the UEs without their involvement or service interruption.
@@
After measuring neighbor cells, we determine if handover to the neighbor gNB improves signal quality for all UEs inside the vehicle. Standard A3 HO triggers when neighbor RSRP ($M_{n,i}$) exceeds serving RSRP ($M_{s,i}$) by a threshold $H$ ($M_{n,i}>M_{s,i}+H$) for $i$-th UE. However, this per-user HO could lead to conflicting decisions for different UEs. Also, the serving gNB measures the reflected neighbor signal ($M_{r}$), not the required $M_{n,i}$. As seen in Fig.˜8, we define the serving, neighbor, and reflected RSRP as:
(4) $M_{s,i}$ $=(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})+G_{\mathrm{w,t}}+(G_{ue,i}+L_{ue,i})$
(5) $M_{n,i}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,t}}+(G_{ue,i}+L_{ue,i})$
(6) $M_{r}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,r}}+(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})$
The serving gNB directly measures $M_{s,i}$ from each UE and $M_{r}$ from Wall-Street. It also knows surface gains $G_{w,t}$ and $G_{w}$ from the codebook and the UE gain $G_{ue,i}$ [^4] In 3GPP, the gNB sends Downlink Control Information (DCI) to send power control commands to the UE.. The UE-specific path loss $L_{ue,i}$ remains unknown.
Instead of using gNB-UE links, our approach is to make one HO decision for the entire vehicle based on the predictable gNB-surface links. We redefine the serving link as $X_{s}=G_{gNB,s}+L_{gNB,s}$ and the neighbor link as $X_{n}=G_{gNB,n}+L_{gNB,n}$. Then, the HO decision for all $K$ users becomes $X_{n}>X_{s}+H$. Our goal is to estimate $X_{s}$ and $X_{n}$ from available measurements. To do so, we employ a two-slot measurement with bounding algorithm described next.
Figure 8. Measurement power translation: Inferring 3GPP-standard measurement report power readings based on measurements reflected from Wall-Street.
Two-slot measurements. We use Wall-Street’s dynamic power allocation to estimate $X_{s}$ and $X_{n}$ accurately. It allocates power between reflection ($\alpha_{j}$) and refraction ($1-\alpha_{j}$) in each time slot $j$. In Slot 1, we increase reflection to measure accurate $M_{r}$ (already measured in Section˜4.2.2), while in Slot 2, we increase refraction to accurately measure $M_{s,i}$. Based on Eq.˜4 and Eq.˜6, we define the per-slot models as:
(7) $M_{r}^{(j)}=X_{n}+X_{s}+G_{w,r}(\alpha_{j})$
(8) $M_{s,i}^{(j)}=X_{s}+G_{w,t}(1-\alpha_{j})+G_{ue,i}+L_{ue,i}^{(j)}$
@@
Based on Eq.˜9, we can write our decision metric as $\Delta=X_{n}-X_{s}=S_{1}-2X_{s}$, so the bounds for our metric becomes $\Delta_{min}=S_{1}-2UBs$. We trigger HO if $\Delta_{min}\geq H$. Otherwise, HO does not occur to avoid unnecessary HOs. This conservative approach prevents rapid ping-pong HOs caused by transient signal fluctuations.
@@ 4.2.4. Handover Execution
The source gNB initiates the HO by sending a HO Req message to the target gNB. Upon receiving the HO ACK, the source gNB triggers Wall-Street to enter the make-before-break mode and sends HO command to the UE. The UE synchronizes with the target gNB via RACH and maintains two active protocol stacks using the 3GPP-standardized Dual Active Protocol Stack (DAPS) (3GPP, 2021). It establishes new Signaling Radio Bearers (SRBs) for control-plane signaling (including Wall-Street control) with the target cell while suspending the source cell SRBs. For the user plane, a common Packet Data Convergence Protocol (PDCP) entity is created to serve both links, allowing the UE to exchange data with both gNBs on different channels.
At the physical layer, Wall-Street combines downlink beams from both gNBs and splits the uplink beam towards both gNB. No additional beam searching is required for this step as the optimal angles for the serving link ($\theta_{1}$) and the target link ($\theta_{2}$) are already known from HO preparation. To ensure reliable packet delivery, both the source and target gNBs transmit duplicate data packets to the UE, where the PDCP layer handles reordering and de-duplication.
Wall-Street initially divides power equally between the two links. If one link degrades, it temporarily boosts that path to maintain a minimum quality of service, otherwise it falls back to an even split. During make-before-break, Wall-Street can directly measure both $M_{s}$ and $M_{n}$ to check A3 condition in real time. If the condition is not met for a TTT duration, the system reverts to the serving cell.
@@ 4.2.5. Handover Completion
The target cell completes the HO by sending a UE release to the source cell and HO completion to Wall-Street. Upon completion, Wall-Street maximizes transmission power to the new gNB. If the HO fails, Wall-Street instead maximizes power to the old gNB so that UE does not restart from the initial attachment phase. The UE also seamlessly reverts to the source gNB configuration and reactivates source cell SRBs for control-plane signaling. We note that a single HO execution simultaneously transfers all UEs within the same vehicle, reducing control-plane signaling.
@@
Figure 11. Experimental gNB-side testbed setup and map: we deploy gNBs on the first floor of a lab facing an outdoor parking lot. The vehicle is approximately 2020 to $30$ m from the gNBs. The arrow indicates the direction that each gNB node is facing at 0°, and each node sweeps the beam from -60° to 60°.
5. Implementation
Figure˜9 illustrates our implementation of Wall-Street. It consists of (1) a surface hardware and control unit; (2) integration with the COSMOS testbed, implementing new PHY features; and (3) an ns-3 simulation for evaluating TCP end-to-end performance, using a state-of-the-art mmWave ns-3 simulator (Mezzavilla et al., 2018) with our own modifications. The simulation is trace-driven, fed with PHY traces collected from COSMOS testbed in real-time under a multi-user scenario.
5.1. Wall-Street Hardware
The Wall-Street surface is composed of 76 boards, each consisting of 28 co-located magnetic and electric unit cells as shown in Fig.˜10(a). Two 40-channel AD5370 16-bit DACs provide independent control of the unit cells on each Wall-Street board, with each channel supplying a variable 0 to 16 V control voltage. To expedite the steering process, one channel supplies the voltage to two adjacent boards, enabling two DACs to independently control all boards. The two DACs are each connected to the Serial Peripheral Interface (SPI) of a Raspberry Pi (RPI) through GPIO. The RPI listens for control signals from the COSMOS testbed control network, known as sandbox1 (SB1), via sub-6 GHz. Upon receiving a signal, it sends a command to the DACs, which then apply the appropriate voltage levels. These voltage levels are determined from a pre-stored angle-to-voltage codebook. For multi-beams, we construct codebook for different power ratios between two beams ($\alpha=0.25,0.5,0.75$). The speed of the Wall-Street hardware has been optimized to 0.2 ms for real-time experiments.
@@ 5.2. COSMOS Testbed
The COSMOS testbed comprises software-defined radios (USRPs) integrated with the IBM mmWave Phased-Array Antenna Module (PAAM) frontends (Fig.˜9). The PAAM features 4 tiled RFICs, each with 32 TRx phase-shifting elements, supporting a total of 64 antennas. The PAAM offers full access to beamforming control and latency as low as 10s of nanoseconds for low-latency MAC and hybrid beamforming. As shown in Fig.˜10(a), our experimental setup uses two types of PAAM devices: two stationary PAAM nodes integrated with USRP N310s mounted on remotely controllable XY tables, and three portable PAAM nodes integrated with USRP 2974s. For indoor scenarios, we utilize four gNBs: two stationary PAAMs (gNB 1 and 4) with a 35 dB gain and two portable PAAMs (gNB 2 and 3) with a 31 dB gain. In multi-UE outdoor scenarios, we use three gNBs: two stationary PAAMs (gNB 1 and 2) with a 65 dB gain and one portable PAAM (gNB 3) with a 31 dB gain, with the remaining portable PAAMs serving as UEs with a 31 dB gain [^5] The node gains are defined by the COSMOS testbed. We refer the reader to https://www.cosmos-lab.org/technology/mmwave/ for the details.. The location of UEs remain the same with and without the surface. Since all nodes must be connected to the same control network (SB1), we use a Wi-Fi bridge operating at sub-6 GHz to send the control commands to outdoor UEs and the surface.
@@ 5.2.1. PHY implementation on COSMOS
We developed new features on the PHY layer of the COSMOS testbed, including OFDM packet transmission and reception on two different channels, extraction of RSRP and Packet Error Rate (PER) from the packet, and IBM PAAM and Wall-Street beam control. Periodic bursts of OFDM packets are sent from two PAAMs in different frequency bands with 3.84 MHz separation, and the receiver simultaneously decodes packets from both transmitters by operating at a larger bandwidth. Specifically, the receiver extracts RSRP from the reference signal in the preamble, the packet sequence number from the header, and the PER from the cyclic redundancy check (CRC). During the make-before-break handover, the source and target gNBs transmit identical packets, and the UE decodes packets on both channels simultaneously and compute the combined PER. A C++ script in SB1 sends control commands to the gNB and Wall-Street to follow the handover procedure, as illustrated in Fig.˜9 (right).
Figure 12. Beam patterns for four steering modes. Red/blue $\triangle$ show measurements at 26.0/26.1 GHz. $\blacksquare$ show baseline (surface turned off). Gray lines indicate target angles. Left: -40°/40° steering; right: -45°/68° steering.

### v4

@@ 4.2. Wall-Street Handover Design
This section describes Wall-Street’s HO process, which is divided into: initial attachment, preparation, decision, execution, and completion, as detailed in Fig.˜4. First, to ensure uninterrupted connectivity, we offload the neighbor cell measurement task from users to the surface (Section˜4.2.2). Second, to improve efficiency, we batch the HO for all in-vehicle users into a single operation (Section˜4.2.3). Finally, to maximize reliability, we establish a make-before-break connection, linking users to the new cell before disconnecting from the old cell (Section˜4.2.4).
System Assumptions. Our design operates under three architectural assumptions:
  • $\bullet$
    Nearby gNBs transmit their broadcast signals on different frequency resources. This is a standard deployment practice to prevent signal collisions. We leverage this practice for both handover preparation and execution.
  • $\bullet$
    The serving gNB can simultaneously measure reflected signals from a neighbor while maintaining communication with the UE on a separate frequency.
  • $\bullet$
    The serving gNB controls Wall-Street via a sub-6 GHz control channel. Instead of accessing the surface’s low-level codebooks, gNBs signals high-level beam steering angles. The surface’s onboard microcontroller has the hardware-specific codebook and locally maps angles into the voltage configurations.
We further discuss protocol integration, 3GPP compatibility, and methods to reduce system complexity in Section˜7.
@@ 4.2.1. Initial Attachment
As illustrated in Fig.˜5(a), the initial attachment requires a coordinated beam sweep to align the gNB, Wall-Street (operating in a single-beam transmission mode), and the UE. Every $20$ ms period, the gNB broadcasts a $5$ ms burst containing 64 synchronization signal blocks (SSBs), which are transmitted across 8 different gNB Tx beam directions. For each SSB, Wall-Street refract them into $8$ angles, testing different paths into the vehicle. For each burst, the UE holds its Rx beam and measures the power. The UE tests a different beam direction for each burst. After four bursts (one per UE Rx direction) in $80$ ms, the UE reports the best alignment and connects with the gNB.
@@ 4.2.2. Handover Preparation
The serving gNB periodically triggers Wall-Street to enter a dual-beam mode for neighbor cell measurement. The goal is to measure neighbor cell RSRP without interrupting data sessions, as shown in Fig.˜4(b). Since the refractive link between the serving gNB, Wall-Street, and the UE is already established, the transmissive angle $\theta_{t}$ is known. Wall-Street then uses its pre-computed codebook indexed by $(\theta_{t},\theta_{r},\alpha)$ to cycle through 8 reflective angles $\theta_{r}$, and neighbor gNBs transmit SSBs in their standard beam patterns. Since the serving gNB already knows its direction toward Wall-Street, it measures these reflected SSBs on a separate channel to prevent interference with the active data link. The power split $\alpha$ allocates less power to maintain the data connection while assigning the remaining power to the reflective link. As illustrated in Fig.˜5(b), this search takes one 5 ms burst. This allows the serving gNB to measure the neighbor RSRPs on behalf of the UEs without their involvement or service interruption.
@@
After measuring neighbor cells, the serving gNB determines if new handover will improve signal quality for all UEs inside the vehicle. Standard A3 HO triggers when the neighbor RSRP ($M_{n,i}$) exceeds the serving RSRP ($M_{s,i}$) by a threshold $H$ ($M_{n,i}>M_{s,i}+H$) for $i$-th UE. However, this per-user HO could lead to conflicting decisions for different UEs. Also, the serving gNB measures the reflected signal $M_{r}$ instead of $M_{n,i}$. As illustrated in Fig.˜6, we simplify the serving, neighbor, and reflected RSRP as:
(2) $M_{s,i}$ $=(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})+G_{\mathrm{w,tra}}+(G_{ue,i}+L_{ue,i})$
(3) $M_{n,i}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,tra}}+(G_{ue,i}+L_{ue,i})$
(4) $M_{r}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,ref}}+(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})$
where $G_{gNB,s}$ and $G_{gNB,n}$ denote the serving gNB and neighbor gNB gain, respectively. Similarly, $L_{gNB,s}$ and $L_{gNB,n}$ represent the path loss between the surface and the serving and neighbor gNBs. The serving gNB directly measures $M_{s,i}$ from each UE and $M_{r}$ from Wall-Street. It also knows surface gains $G_{w,tra}$ and $G_{w,ref}$ from the codebook and the UE gain $G_{ue,i}$. [^6] In 3GPP, the gNB sends Downlink Control Information (DCI) to send power control commands to the UE. The UE-specific path loss $L_{ue,i}$ remains unknown.
Instead of using gNB-UE links, our approach is to make one HO decision for the entire vehicle based on the estimated gNB-surface links. We redefine the serving link as $X_{s}=G_{gNB,s}+L_{gNB,s}$ and the neighbor link as $X_{n}=G_{gNB,n}+L_{gNB,n}$. Then, the HO decision for all $K$ users becomes $X_{n}>X_{s}+H$. Our goal is to estimate $X_{s}$ and $X_{n}$ from available measurements. To do so, we employ a two-slot measurement with bounding algorithm described next.
Two-slot measurements. We use Wall-Street’s dynamic power allocation to estimate $X_{s}$ and $X_{n}$ accurately. It allocates power between reflection ($\alpha_{j}$) and refraction ($1-\alpha_{j}$) in each time slot $j$. In Slot 1, we increase reflection to measure accurate $M_{r}$ (already measured in Section˜4.2.2), while in Slot 2, we increase refraction to accurately measure $M_{s,i}$. Based on Eq.˜2 and Eq.˜4, we define the per-slot models as:
(5) $M_{r}^{(j)}=X_{n}+X_{s}+G_{w,ref}(\alpha_{j})$
(6) $M_{s,i}^{(j)}=X_{s}+G_{w,tra}(1-\alpha_{j})+G_{ue,i}+L_{ue,i}^{(j)}$
@@
Based on Eq.˜7, we can write our decision metric as $\Delta=X_{n}-X_{s}=S_{1}-2X_{s}$, so the bounds for our metric becomes $\Delta_{min}=S_{1}-2UBs$. We trigger HO if $\Delta_{min}\geq H$. Otherwise, HO does not occur to avoid unnecessary HOs. This conservative approach prevents rapid ping-pong HOs caused by transient signal fluctuations.
Figure 6. Measurement power translation: Inferring measurement report power readings based on measurements reflected from Wall-Street.
@@ 4.2.4. Handover Execution
The serving gNB initiates the HO by sending a HO Req message to the target gNB. Upon receiving the HO ACK, the serving gNB triggers Wall-Street to enter a make-before-break mode and sends HO command to UEs. UEs synchronize with the target gNB via RACH and maintain two active protocol stacks using the 3GPP-standardized Dual Active Protocol Stack (DAPS) (3GPP, 2021a). It establishes new Signaling Radio Bearers (SRBs) for control-plane signaling, including the Wall-Street control, with the target cell while suspending the serving cell SRBs. For the user plane, a common Packet Data Convergence Protocol (PDCP) entity is created to serve both links, allowing UEs to exchange data with both gNBs on different channels.
At the physical layer, Wall-Street combines downlink beams from both gNBs and splits the uplink beam towards both gNBs. No additional beam searching is required for this step as the optimal angles for the serving link ($\theta_{1}$) and the target link ($\theta_{2}$) are already known from HO preparation. To ensure reliable packet delivery, both the serving and target gNBs transmit duplicate data packets to UEs, where the PDCP layer handles reordering and de-duplication. During make-before-break, Wall-Street can directly measure both $M_{s}$ and $M_{n}$ to check A3 condition in real time. A3 conditions are continuously evaluated. If the condition is not met, the system reverts to the serving cell.
@@ 4.2.5. Handover Completion
The target cell completes the HO by sending a UE release to the serving cell and HO completion to Wall-Street. Upon completion, Wall-Street maximizes transmission power to the new gNB. If the HO fails, Wall-Street instead maximizes power to the old gNB so that UEs do not restart from the initial attachment phase. UEs also reverts to the old gNB configuration and reactivates old cell SRBs for control-plane signaling. We note that a single HO execution simultaneously transfers all UEs within the same vehicle, reducing control-plane signaling.
5. Implementation
Figure˜7 illustrates our implementation of Wall-Street. It consists of (1) a surface hardware and control unit; (2) integration with the COSMOS testbed, implementing new PHY features; and (3) an mmWave ns-3 simulation (Polese, 2019) for evaluating TCP performance, which was modified to be trace-driven and fed with PHY traces collected from COSMOS testbed.
5.1. Wall-Street Hardware
The Wall-Street surface is composed of 76 boards, each consisting of 28 co-located magnetic and electric unit cells as shown in Fig.˜8(a). Two 40-channel AD5370 16-bit DACs provide independent control of the unit cells on each board, with each channel supplying a variable 0 to 16 V control voltage. To expedite the steering process, one channel supplies the voltage to two adjacent boards, enabling two DACs to independently control all boards. The two DACs are each connected to the Serial Peripheral Interface (SPI) of a Raspberry Pi (RPI) through GPIO. The RPI listens for control signals from the COSMOS testbed control network, known as sandbox1 (SB1), via sub-6 GHz. Upon receiving a signal, it sends a command to the DACs, which then apply the appropriate voltage levels. These voltage levels are determined from a pre-stored angle-to-voltage codebook. For multi-beams, we construct codebook for different power ratios between two beams ($\alpha=0.25,0.5,0.75$). The speed of the Wall-Street hardware has been optimized to 0.2 ms for real-time experiments.
@@ 5.1. Wall-Street Hardware
Figure 9. Testbed setup and map: we deploy the gNBs on the first floor of a lab facing an outdoor parking lot. The vehicle is approximately 2020 to $30$ m from the gNBs. The arrow indicates the gNB direction facing at 0°, and each node sweeps the beam from -60° to 60°.
@@ 5.2. COSMOS Testbed
The COSMOS testbed comprises software-defined radios (USRPs) integrated with the IBM mmWave Phased-Array Antenna Module (PAAM) frontends (Fig.˜7). The PAAM features 4 tiled RFICs, each with 32 TRx phase-shifting elements, supporting a total of 64 antennas. The PAAM offers full access to beamforming control and latency as low as 10s of nanoseconds for low-latency MAC and hybrid beamforming. As shown in Fig.˜8(a), our experimental setup has two types of PAAM devices: two stationary PAAM nodes integrated with USRP N310s mounted on remotely controllable XY tables, and three portable PAAM nodes integrated with USRP 2974s. For indoor scenarios, we utilize four gNBs: two stationary PAAMs (gNB 1 and 4) with a 35 dB gain and two portable PAAMs (gNB 2 and 3) with a 31 dB gain. In multi-UE outdoor scenarios, we use three gNBs: two stationary PAAMs (gNB 1 and 2) with a 65 dB gain and one portable PAAM (gNB 3) with a 31 dB gain, with the remaining portable PAAMs serving as UEs [^7] The node gains are defined by the COSMOS testbed. We refer the reader to https://www.cosmos-lab.org/technology/mmwave/ for the details.. The location of UEs remain the same with and without Wall-Street. Since all nodes must be connected to the same control network (SB1), we use a Wi-Fi bridge operating at sub-6 GHz to send the control commands to outdoor UEs and the surface.
@@ 5.2.1. PHY implementation on COSMOS
We developed new features on the PHY layer of the COSMOS testbed, including OFDM packet transmission and reception on two different channels, extraction of RSRP and Packet Error Rate (PER) from the packet, and IBM PAAM and Wall-Street beam control. Periodic bursts of OFDM packets are sent from two PAAMs in different frequency bands with 3.84 MHz separation, and the receiver simultaneously decodes packets from both transmitters by operating at a larger bandwidth. The receiver extracts RSRP from the reference signal in the preamble, the packet sequence number from the header, and the PER from the cyclic redundancy check (CRC). During make-before-break, the serving and target gNBs transmit identical packets, and the UE decodes packets on both channels simultaneously and compute the combined PER. A C++ script in SB1 sends control commands to the gNB and Wall-Street to follow the handover procedure (Fig.˜7 on right).
5.2.2. NS-3 simulation
The COSMOS testbed currently supports mmWave PHY-layer experiments but lacks the full upper-layer stack for TCP traffic. We modified an existing mmWave NS-3 simulator to be trace-driven and fed with the PHY-layer COSMOS traces for TCP analysis. Specifically, we modified its standalone Hard Handover (HH) scheme (we refer to (Polese, 2019) for details). We made major modifications to its PHY and RRC modules to support trace-based operations and handover logic. The PHY module runs a continuous data transmission cycle. Throughout this cycle, the spectrum PHY module continuously updates the data SINR, syncing with timestamps and SINRs from COSMOS data transmission traces. Running independently from the data transmission cycle, a scanning cycle is triggered at a defined measurement report periodicity. During this cycle, the PHY module loops through beam combinations in a 5-ms burst, retrieving SINR values from COSMOS beam scanning traces. Once scanning is complete, the results are reported to the RRC module to evaluate A3 event conditions. The serving cell’s RRC module initiates handover for all UEs when $\Delta_{min}\geq H$.